### v1

@@ 5.1 Validation on synthetic data
Here, we generate each observation $\mathcal{X}$ in $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ by combining a low-rank tensor $\mathcal{L}_{0}$ and a sparse tensor $\mathcal{S}_{0}$ with a Gaussian noise $\mathcal{E}_{0}$ in the the same dimensions. The low-rank tensor $\mathcal{L}_{0}$ is derived from the t-product of two smaller tensors, namely $\mathcal{P}$ in $\mathbb{R}^{n_{1}\times r\times n_{3}}$ and $\mathcal{Q}$ in $\mathbb{R}^{r\times n_{2}\times n_{3}}$, where $r$ is significantly smaller than $n_{2}$. The tubal rank of $\mathcal{L}_{0}$ does not exceed $r$. The entries of $\mathcal{P}$ are independently and identically distributed according to a Gaussian distribution $\mathcal{N}(0,1/n_{1})$, and those of $\mathcal{Q}$ follow $\mathcal{N}(0,1/n_{2})$. The sparse tensor $\mathcal{S}_{0}$ has entries determined by a Bernoulli process, where each element is either $+1$ or $-1$ with a probability $\rho$, and 00 with a probability $1-2\rho$. The entries in Gaussian noise $\mathcal{S}_{0}$ follow $\mathcal{N}(0,\sigma^{2})$.
@@
We initiate our analysis by examining the convergence properties using a third-order tensor with dimensions $40\times 40\times 30$. The rank parameter $r$ is set to 3, with the regularization parameter $\rho$ at 0.1 and the noise level $\sigma$ at $10^{-2}$. The algorithm is allowed a maximum of 100 iterations, starting with initial guesses for $\mathcal{L}$ and $\mathcal{S}$ as $\mathcal{X}$ and $\mathcal{O}$, respectively. The convergence of the algorithm is monitored using the relative mean square error (RMSE) for $\mathcal{L}$ and $\mathcal{S}$, defined as $\frac{\|\mathsf{E}_{\mathcal{L}}^{\ell}-\mathsf{E}_{\mathcal{L}}^{\ell-1}\|_{F}}{\|\mathsf{E}_{\mathcal{L}}^{\ell}\|_{F}}$ and $\frac{\|\mathsf{E}_{\mathcal{S}}^{\ell}-\mathsf{E}_{\mathcal{S}}^{\ell-1}\|_{F}}{\|\mathsf{E}_{\mathcal{S}}^{\ell}\|_{F}}$, respectively. The progression of the objective values, RMSE, and parameters ($\theta_{1},\theta_{2},\theta_{3}$) are plotted across iterations in Figure 1. Due to the nonlinear and nonconvex nature of simultaneously optimizing three tensors and their associated parameters, initial fluctuations in the objective values are observed. However, after approximately ten iterations, the objective values begin to decrease steadily and achieve convergence by the 30th iteration. The parameter values similarly stabilize within these iterations. Both RMSE metrics show a sharp decline, reaching as low as $10^{-4}$ by the 30th iteration. Given these observations, we establish a stopping criterion where the algorithm terminates when RMSE falls below $10^{-4}$ or when 50 iterations are reached, whichever occurs first. This criterion ensures efficient and effective convergence to an optimal solution within a reasonable number of iterations.
@@
As part of a proof-of-concept study, we employ a partial sum of the tubal nuclear norm [22] as a representative example for a weighted TNN in our numerical experiments. We aim to compare our proposed algorithms, VBITNN and VBIPSTNN, against two established methods in tensor rank approximation: TNN [32] and PSTNN [22]. For this comparative analysis, we set the noise levels $\sigma$ at $10^{-3}$, $10^{-2}$, and $10^{-1}$, the rank $r$ at 3 and 5, and the regularization parameter $\rho$ at 0.01 and 0.1. We assess the performance of these methods by calculating the relative square error between the recovered tensors, $\hat{\mathcal{L}}$ and $\hat{\mathcal{S}}$, and the ground-truth tensors, $\mathcal{L}_{\rm GT}$ and $\mathcal{S}_{\rm GT}$. These errors are quantified as follows: ${\rm error}_{\mathcal{L}}=\frac{\|\hat{\mathcal{L}}-\mathcal{L}_{\rm GT}\|_{F}}{\|\mathcal{L}_{\rm GT}\|_{F}}$ for the low-rank component and ${\rm error}_{\mathcal{S}}=\frac{\|\hat{\mathcal{S}}-\mathcal{S}_{\rm GT}\|_{F}}{\|\mathcal{S}_{\rm GT}\|_{F}}$ for the sparse component.
@@
As shown in Table 1, VBITNN generally outperforms TNN across most tested scenarios, while VBIPSTNN is better than PSTNN. Moreover, VBIPSTNN consistently delivers the best performance, indicating its superior ability to recover both the low-rank and sparse components of tensors under various noise and rank conditions. This comparative analysis underscores the effectiveness of our proposed methods, particularly VBIPSTNN, in handling complex tensor decomposition with higher accuracy and robustness against noise.
@@
Figure 1: Empirical evidence on convergence. Left: objective function, middle: RMSE, right: parameters: $\theta_{1},\theta_{2}$, and $\theta_{3}$, generated by Algorithm 1 across iterations.
@@ 5.2.1 Image with sparse noise
We conduct experiments on four images: “house,” “moto,” “face,” and “hat”. In this study, we model the clean images as the low-rank component and random corruptions as sparse outliers. Each image is corrupted by setting 10 percent of the pixels to random values ranging from 0 to 255, with the locations of these distortions unspecified. We compare our proposed method with several existing techniques, including LRTV [19], $S_{wp}(0.9)$ [49], BTRTF [53], TNN [32], and PSTNN [22], using the original implementations provided by the respective authors. Given the absence of Gaussian noise in this task, the parameter $\theta_{1}$ is set to a high value of 100 to accommodate this condition, while $\theta_{2}$ and $\theta_{3}$ are set to 1. The truncation parameter $K$ for VBIPSTNN is consistently set at 50 across all cases.
Quantitative evaluations based on PSNR and SSIM are presented in Table 2, and the corresponding restored images are displayed in Figure 2. Our observations indicate that VBIPSTNN consistently outperforms the other methods in terms of PSNR, achieving at least a 0.5 improvement and matching the best-performing methods in SSIM values. Additionally, the restoration of the “hat” image by VBIPSTNN and BTRTF shows significantly clearer text compared to other methods. However, some artifacts are noted in the “moto” image restored by BTRTF. In contrast, our method exhibits fewer artifacts across all cases.
@@ 5.2.2 Image with mixed noise
We benchmark our proposed algorithm against several state-of-the-art methods, including 3DTNN [51], $S_{wp}(0.9)$ [49], BTRTF [53], TNN [32], and PSTNN [22]. Performance metrics such as PSNR and SSIM are detailed in Table 3, with visual results presented in Figure 3. Notably, our algorithm outperforms both TNN and PSTNN—methods that utilize similar regularization techniques—across all test cases in terms of PSNR, achieving an average improvement of 0.6 dB over the best-reported results. Qualitatively, the images restored by VBIPSTNN exhibit notably sharper boundaries compared to those produced by the other methods, which tend to exhibit some degree of blurring.
@@ 5.3 Background modeling
We evaluated our models on sequences from the 12R dataset [31], specifically the “bootstrap” ($120\times 160\times 400$), and “sidewalk” ($220\times 352\times 400$) videos, all characterized by slow-moving objects against varying backgrounds. Our models were compared with several others, including 3DTNN, TNN, BTRTF, PSTNN, and $\mathrm{t}\mbox{-}S_{w,p}(0.9)$. For VBIPSTNN, the truncated parameter $K$ is set as 5, while the initial values of $\theta_{1},\theta_{2},\theta_{3}$ are set as 1, 1, 100, respectively. The results of these comparisons are visually presented in Figure 4. Each video’s analysis starts with a frame from the sequence as shown in column (a) of Figure 4, followed by background images generated by the respective methods, from 3DTNN to our approach VBIPSTNN. Additionally, the motion in each scene is depicted in the second row for each video.
In the “bootstrap” video, except for 3DTNN, all the methods achieved superior background separation with fewer ghost silhouettes. When we zoomed in on the images, we observed that our approach has a background in $\mathcal{S}_{0}$ than the others, which leads to better separation. In the “sidewalk” videos, all the approaches perform similarly, while 3DTNN has slightly better results.

### v2

@@ 5.1 Validation on synthetic data
Here, we generate each observation $\mathcal{X}$ in $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ by combining a low-rank tensor $\mathcal{L}_{0}$ and a sparse tensor $\mathcal{S}_{0}$ with a Gaussian noise $\mathcal{E}_{0}$ in the the same dimensions. The low-rank tensor $\mathcal{L}_{0}$ is derived from the t-product of two smaller tensors, namely $\mathcal{P}$ in $\mathbb{R}^{n_{1}\times r\times n_{3}}$ and $\mathcal{H}$ in $\mathbb{R}^{r\times n_{2}\times n_{3}}$, where $r$ is significantly smaller than $n_{2}$. The tubal rank of $\mathcal{L}_{0}$ does not exceed $r$. The entries of $\mathcal{P}$ are independently and identically distributed according to a Gaussian distribution $\mathcal{N}(0,1/n_{1})$, and those of $\mathcal{H}$ follow $\mathcal{N}(0,1/n_{2})$. The sparse tensor $\mathcal{S}_{0}$ has entries determined by a Bernoulli process, where each element is either +1+1 or 1-1 with a probability $\rho$, and 0 with a probability $1-2\rho$. The entries in Gaussian noise $\mathcal{S}_{0}$ follow $\mathcal{N}(0,\sigma^{2})$.
@@
We initiate our analysis by examining the convergence properties using a third-order tensor with dimensions $40\times 40\times 30$. The rank parameter $r$ is set to 3, with the parameter $\rho$ at 0.1 and the noise level $\sigma$ at $10^{-2}$. The algorithm is allowed a maximum of 100 iterations, starting with initial guesses for $\mathcal{L}$ and $\mathcal{S}$ as $\mathcal{X}$ and $\mathcal{O}$, respectively. The convergence of the algorithm is monitored using the relative mean square error (RMSE) for $\mathcal{L}$ and $\mathcal{S}$, defined as $\frac{\|\mathsf{E}_{\mathcal{L}}^{\ell}-\mathsf{E}_{\mathcal{L}}^{\ell-1}\|_{F}}{\|\mathsf{E}_{\mathcal{L}}^{\ell}\|_{F}}$ and $\frac{\|\mathsf{E}_{\mathcal{S}}^{\ell}-\mathsf{E}_{\mathcal{S}}^{\ell-1}\|_{F}}{\|\mathsf{E}_{\mathcal{S}}^{\ell}\|_{F}}$, respectively. The progression of the objective values, RMSE, and parameters ($\theta_{1},\theta_{2},\theta_{3}$) is plotted across iterations in (1). Due to the nonlinear and nonconvex nature of simultaneously optimizing three tensors and their associated parameters, initial fluctuations in the objective values are observed. However, after approximately ten iterations, the objective values begin to decrease steadily and achieve convergence by the 30th iteration. The parameter values similarly stabilize within these iterations. Both RMSE metrics show a sharp decline, reaching as low as $10^{-4}$ by the 30th iteration. Given these observations, we establish a stopping criterion where the algorithm terminates when RMSE falls below $10^{-4}$ or when 50 iterations are reached, whichever occurs first. This criterion ensures efficient and effective convergence to an optimal solution within a reasonable number of iterations.
Here, we further evaluate the uncertainty quantification performance of our Variational Bayesian Inference (VBI) algorithm using the same simulated tensor as previously described. (2). presents the mean estimates and 99.73% credible intervals for the recovery of tensor filter $\overline{\mathcal{L}}_{:ij}$ with $i=20,j=5,15,20$. The mean values consistently align with the ground truth across all fibers, while remarkably narrow credible intervals (indicated by minimal shading) demonstrate the high precision of our method. This precision is further corroborated by the low parameter standard deviations.
@@
As part of a proof-of-concept study, we employ a partial sum of the tubal nuclear norm [23] as a representative example for a weighted TNN in our numerical experiments. We aim to compare our proposed algorithms, VBITNN and VBIPSTNN, against two established methods in tensor rank approximation: TNN [33] and PSTNN [23]. For this comparative analysis, we set the noise levels $\sigma$ at $10^{-3}$, $10^{-2}$, and $10^{-1}$, the rank $r$ at 3 and 5, and the parameter $\rho$ at 0.01 and 0.1. We assess the performance of these methods by calculating the relative square error between the recovered tensors, $\hat{\mathcal{L}}$ and $\hat{\mathcal{S}}$, and the ground-truth tensors, $\mathcal{L}_{\rm GT}$ and $\mathcal{S}_{\rm GT}$. These errors are quantified as follows: ${\rm error}_{\mathcal{L}}=\frac{\|\hat{\mathcal{L}}-\mathcal{L}_{\rm GT}\|_{F}}{\|\mathcal{L}_{\rm GT}\|_{F}}$ for the low-rank component and ${\rm error}_{\mathcal{S}}=\frac{\|\hat{\mathcal{S}}-\mathcal{S}_{\rm GT}\|_{F}}{\|\mathcal{S}_{\rm GT}\|_{F}}$ for the sparse component.
@@
As shown in (1), VBITNN generally outperforms TNN across most tested scenarios, while VBIPSTNN is better than PSTNN. Moreover, VBIPSTNN consistently delivers the best performance, indicating its superior ability to recover both the low-rank and sparse components of tensors under various noise and rank conditions. This comparative analysis underscores the effectiveness of our proposed methods, particularly VBIPSTNN, in handling complex tensor decomposition with higher accuracy and robustness against noise.
@@
Figure 1: Empirical evidence on convergence. Left: objective function, middle: RMSE, right: parameters: $\theta_{1},\theta_{2}$, and $\theta_{3}$, generated by (1) across iterations.
Refer to caption Refer to caption Refer to caption
Figure 2: Uncertainty quantification: recovery of $\overline{\mathcal{L}}_{:ij}$ with 99.73% credible interval (shaded area) where $i=20,j=5,10,25$.
@@ 5.2.1 Image with sparse noise
We conduct experiments on four images: “house”, “moto”, “face”, and “hat”. In this study, we model the clean images as the low-rank component and random corruptions as sparse outliers. Each image is corrupted by setting 10 percent of the pixels to random values ranging from 0 to 255, with the locations of these distortions unspecified. We compare our proposed method with several existing techniques, including LRTV [20], $S_{wp}(0.9)$ [51], BTRTF [55], TNN [33], and PSTNN [23], using the original implementations provided by the respective authors. Given the absence of Gaussian noise in this task, the parameter $\theta_{1}$ is set to a high value of 100 to accommodate this condition, while $\theta_{2}$ and $\theta_{3}$ are set to 1. The truncation parameter $K$ for VBIPSTNN is consistently set at 50 across all cases.
Quantitative evaluations based on PSNR and SSIM are presented in (2), and the corresponding restored images are displayed in (3). Our observations indicate that VBIPSTNN consistently outperforms the other methods in terms of PSNR, achieving at least a 0.5 improvement and matching the best-performing methods in SSIM values. Additionally, the restoration of the “hat” image by VBIPSTNN and BTRTF shows significantly clearer text compared to other methods. However, some artifacts are noted in the “moto” image restored by BTRTF. In contrast, our method exhibits fewer artifacts across all cases.
@@ 5.2.2 Image with mixed noise
We benchmark our proposed algorithm against several state-of-the-art methods, including 3DTNN [53], $S_{wp}(0.9)$ [51], BTRTF [55], TNN [33], and PSTNN [23]. Performance metrics such as PSNR and SSIM are detailed in (3), with visual results presented in (4). Notably, our algorithm outperforms both TNN and PSTNN—methods that utilize similar regularization techniques—across all test cases in terms of PSNR, achieving an average improvement of 0.6 dB over the best-reported results. Qualitatively, the images restored by VBIPSTNN exhibit notably sharper boundaries compared to those produced by the other methods, which tend to exhibit some degree of blurring.
@@ 5.3 Background modeling
We evaluated our models on sequences from the 12R dataset [32], specifically the “bootstrap” ($120\times 160\times 400$), and “sidewalk” ($220\times 352\times 400$) videos, all characterized by slow-moving objects against varying backgrounds. Our models were compared with several others, including 3DTNN, TNN, BTRTF, PSTNN, and $\mathrm{t}\mbox{-}S_{w,p}(0.9)$. For VBIPSTNN, the truncated parameter $K$ is set as 5, while the initial values of $\theta_{1},\theta_{2},\theta_{3}$ are set as 1, 1, 100, respectively. The results of these comparisons are visually presented in Figure 5. Each video’s analysis starts with a frame from the sequence as shown in column (a) of Figure 5, followed by background images generated by the respective methods, from 3DTNN to our approach VBIPSTNN. Additionally, the motion in each scene is depicted in the second row for each video. In the “bootstrap” video, except for 3DTNN, all the methods achieved superior background separation with fewer ghost silhouettes. In the “sidewalk” videos, all the approaches perform similarly, while 3DTNN has slightly better results.